# Second- and third-order properties of multidimensional Langevin equations

### v16

@@ 8.3.3. Quantitative significance
(365) ×$\quad{}\times\left[{L(\widetilde{x}^{i}\widetilde{v}^{j^{\prime}},(\mathbf{A}_{\mathbf{v}}\widetilde{\mathbf{x}})^{k^{\prime}})+L(\widetilde{x}^{i}\widetilde{v}^{k^{\prime}},(\mathbf{A}_{\mathbf{v}}\widetilde{\mathbf{x}})^{j^{\prime}})-\mu^{-1}L(\widetilde{v}^{j^{\prime}}\widetilde{v}^{k^{\prime}},\widetilde{x}^{i})}\right]+\mathcal{O}(\mu^{-2}).$
@@
×$\quad{}\times\langle x^{i}x^{i^{\prime}}\rangle\Big{[}{D^{ll^{\prime}}\left({\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{D}{\mathbf{A}_{\mathbf{v}}^{-1}}^{\mathsf{T}}}\right)^{mm^{\prime}}-\left({\mathbf{D}{\mathbf{A}_{\mathbf{v}}^{-1}}^{\mathsf{T}}}\right)^{lm^{\prime}}(\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{D})^{ml^{\prime}}}\Big{.}$
@@
+.$\qquad{}+\Big{.}{D^{mm^{\prime}}\left({\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{D}{\mathbf{A}_{\mathbf{v}}^{-1}}^{\mathsf{T}}}\right)^{ll^{\prime}}-\left({\mathbf{D}{\mathbf{A}_{\mathbf{v}}^{-1}}^{\mathsf{T}}}\right)^{ml^{\prime}}(\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{D})^{lm^{\prime}}}\Big{]}+\mathcal{O}(\mu^{0}),$
@@ 8.4. Dynamics on long time-scales
$\mathbf{f}_{1\widetilde{\mathbf{x}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}})$ $=\widetilde{\mathbf{x}}+\mathcal{O}(\mu^{-1})a\mathcal{O}(a,b)\widetilde{\mathbf{v}}+\mathcal{O}(a)(\widetilde{x}^{i}\widetilde{x}^{j}-\langle\widetilde{x}^{i}\widetilde{x}^{j}\rangle)$
@@
$\mathbf{f}_{1\widetilde{\mathbf{v}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}})$ $=\widetilde{\mathbf{v}}+a\mathcal{O}(a,b)\widetilde{\mathbf{x}}+\mathcal{O}(\mu^{-1})\mathcal{O}(a)(\widetilde{x}^{i}\widetilde{x}^{j}-\langle\widetilde{x}^{i}\widetilde{x}^{j}\rangle)$
@@
(371) $\left\langle\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{f}_{1\widetilde{\mathbf{x}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}})\mid\widetilde{\mathbf{x}},\widetilde{\mathbf{v}}\right\rangle$ $=-\mu^{-1}\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{A}_{\mathbf{x}}[1+\mathcal{O}(\mu^{-2})+\mathcal{O}(\mu)a\mathcal{O}(a,b)]\mathbf{f}_{1\widetilde{\mathbf{x}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}}),$
(372) $\left\langle\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{f}_{1\widetilde{\mathbf{v}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}})\mid\widetilde{\mathbf{x}},\widetilde{\mathbf{v}}\right\rangle$ $=\mu\mathbf{A}_{\mathbf{v}}[1+\mathcal{O}(\mu^{-2})+a\mathcal{O}(a,b)]\mathbf{f}_{1\widetilde{\mathbf{v}}}(\widetilde{\mathbf{x}},\widetilde{\mathbf{v}}).$
@@ 9. Detection of non-Markovianity
($\tau\geq 0$), where $-\mu_{\pm}$ are the eigenvalues of $\mathbf{A}$:
@@ 12. Appendix B
Applying Itô’s lemma and taking expectations apparently gives $2\langle x^{2}\rangle=-2\langle x\dot{x}\rangle=1+\langle x^{6}\rangle$. However, the l.h.s. is finite whereas the r.h.s. is infinite [^24] The same problem is encountered if $x^{6}$ is replaced by $x^{4}$ in Eq. (402).. The problem lies in the middle expression. Upon Euler–Maruyama discretization with a fixed time-step, i.e., $x_{i+1}=(1-\Delta t)x_{i}+\sqrt{(1+x_{i}^{6})\Delta t}\cdot\zeta_{i}$ with $\zeta_{i}\sim\mathcal{N}(0,1)$ i.i.d., it would seem that $\langle x_{i}(x_{i+1}-x_{i})/\Delta t\rangle$ should be $-\langle x^{2}\rangle$ because $\langle x^{4}\rangle<\infty$. However, with a fixed time-step $\Delta t$, the sequence $x_{i}$ diverges. To properly simulate this system, an adaptive time-step must be used, i.e., $\Delta t_{i}$ depending on $x_{i}$. We simulated this system using $\Delta t_{i}=0.05(1+x_{i}^{4})^{-1}$ for $10^{8}$ time-steps. Numerically, it appears that $\langle x\dot{x}\rangle=-\infty$ (Fig. 7), in accordance with Eq. (5). Moreover, although the covariance function $\langle x(0)x(\tau)\rangle$ exists for all $\tau$, it is apparently not equal to $\langle x^{2}\rangle e^{-\tau}$ (Fig. 8, calculated using $3\times 10^{5}$ time-steps), as might be expected from the drift function. Thus stochastic force inference fails when using a linear basis function. However, binning reproduces the correct drift (Fig. 9, calculated using $10^{7}$ time-steps). We thus conclude that there is a non-commutation of limits:

### v17

@@ 8.3.3. Quantitative significance
(361) $\quad{}\times\left[{L(\widetilde{x}^{i}\widetilde{v}^{j^{\prime}},(\mathbf{A}_{\mathbf{v}}\widetilde{\mathbf{x}})^{k^{\prime}})+L(\widetilde{x}^{i}\widetilde{v}^{k^{\prime}},(\mathbf{A}_{\mathbf{v}}\widetilde{\mathbf{x}})^{j^{\prime}})-\mu^{-1}L(\widetilde{v}^{j^{\prime}}\widetilde{v}^{k^{\prime}},\widetilde{x}^{i})}\right]+\mathcal{O}(\mu^{-2}).$
@@
$\quad{}\times\langle x^{i}x^{i^{\prime}}\rangle\Big[{D^{ll^{\prime}}\left({\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{D}{\mathbf{A}_{\mathbf{v}}^{-1}}^{\mathsf{T}}}\right)^{mm^{\prime}}-\left({\mathbf{D}{\mathbf{A}_{\mathbf{v}}^{-1}}^{\mathsf{T}}}\right)^{lm^{\prime}}(\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{D})^{ml^{\prime}}}\Big$.
@@
$\qquad{}+\Big.{D^{mm^{\prime}}\left({\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{D}{\mathbf{A}_{\mathbf{v}}^{-1}}^{\mathsf{T}}}\right)^{ll^{\prime}}-\left({\mathbf{D}{\mathbf{A}_{\mathbf{v}}^{-1}}^{\mathsf{T}}}\right)^{ml^{\prime}}(\mathbf{A}_{\mathbf{v}}^{-1}\mathbf{D})^{lm^{\prime}}}\Big]+\mathcal{O}(\mu^{0}),$
@@ 8.4. Dynamics on long time-scales
[…]
@@
[…]
@@
[…]
@@ 9. Detection of non-Markovianity
($\tau\geq 0$), where $\mu_{\pm}$ are the eigenvalues of $\mathbf{A}$:
@@ 12. Appendix B
Applying Itô’s lemma and taking expectations apparently gives $2\langle x^{2}\rangle=-2\langle x\dot{x}\rangle=1+\langle x^{6}\rangle$. However, the l.h.s. is finite whereas the r.h.s. is infinite [^24] The same problem is encountered if $x^{6}$ is replaced by $x^{4}$ in Eq. (398).. The problem lies in the middle expression. Upon Euler–Maruyama discretization with a fixed time-step, i.e., $x_{i+1}=(1-\Delta t)x_{i}+\sqrt{(1+x_{i}^{6})\Delta t}\cdot\zeta_{i}$ with $\zeta_{i}\sim\mathcal{N}(0,1)$ i.i.d., it would seem that $\langle x_{i}(x_{i+1}-x_{i})/\Delta t\rangle$ should be $\langle x^{2}\rangle$ because $\langle x^{4}\rangle<\infty$. However, with a fixed time-step $\Delta t$, the sequence $x_{i}$ diverges. To properly simulate this system, an adaptive time-step must be used, i.e., $\Delta t_{i}$ depending on $x_{i}$. We simulated this system using $\Delta t_{i}=0.05(1+x_{i}^{4})^{-1}$ for $10^{8}$ time-steps. Numerically, it appears that $\langle x\dot{x}\rangle=$-$\infty$ (Fig. 7), in accordance with Eq. (5). Moreover, although the covariance function $\langle x(0)x(\tau)\rangle$ exists for all $\tau$, it is apparently not equal to $\langle x^{2}\rangle e^{-\tau}$ (Fig. 8, calculated using $3\times 10^{5}$ time-steps), as might be expected from the drift function. Thus stochastic force inference fails when using a linear basis function. However, binning reproduces the correct drift (Fig. 9, calculated using $10^{7}$ time-steps). We thus conclude that there is a non-commutation of limits: